OpenStudy (anonymous):

Solve for x. 6x -1/5 = 3/1.

pooja195 (pooja195):

add 1/5 to both sides :)

pooja195 (pooja195):

\[\huge~6x -1/5 +1/5= 3/1+1/5\]

pooja195 (pooja195):

You are left with \[\huge~6x = 3/1+1/5\]

OpenStudy (anonymous):

Thank you so much ! You are very helpful!

pooja195 (pooja195):

haha thanks :) did you get the answer?

OpenStudy (anonymous):

I think so .... 4/6 ? Then simplified ??

pooja195 (pooja195):

hmm now quite :( try doing 3+1/5

pooja195 (pooja195):

again...

OpenStudy (anonymous):

3 and 1/5?

pooja195 (pooja195):

not quite ok you get \[\huge~6x=\frac{ 16 }{ 5 }\] next we want to divide \[\huge~6x/6x=\frac{ 16 }{ 5 }/6x\]

pooja195 (pooja195):

You end up with x=8/15

OpenStudy (anonymous):

thanks

pooja195 (pooja195):

Welcome :) do you get the concept though ?

OpenStudy (anonymous):

I think so ;)

pooja195 (pooja195):

lol if you need help feel free to pm or tag me :)

pooja195 (pooja195):

\(\huge\color{orange}{Welcome~~To~~OpenStudy~:)}\) Hey there!!! Since you are new here, read this legendary tutorial for new OpenStudiers!! http://goo.gl/5pp1u0 For an visual tutorial look at this! Its amazing :o https://prezi.com/fs3hqdpcopic/an-unofficial-guide-to-openstudy/ Also consider reading the CoC and the T&C http://openstudy.com/code-of-conduct http://openstudy.com/terms-and-conditions